I need help with graphing the line y=1/2x + 1. When I did it I got an answer of 0.5 = y.
Am I correct?

Answer by rfer(16322)

It is in y intercept form
(0,1)
(-2,0)
y=1/2x+1
0=1/2x+1
1/2x=-1
x=-1/1*2/1
x=-2/1=-2